Vitex, Inc. manufactures a popular consumer product and it has provided the following data excerpts from its standard cost system:

Inputs (1) Standard Quantity or Hours (2)
Standard
Price
or Rate
Standard
Cost
(1) × (2)
Direct materials 2.30 pounds $ 17.10 per pound $ 39.33
Direct labor 1.00 hours $ 15.90 per hour $ 15.90
Variable manufacturing overhead 1.00 hours $ 9.30 per hour $ 9.30
Total standard cost per unit $ 64.53
Total Variances Reported
Standard
Cost*
Price
or Rate
Quantity or
Efficiency
Direct materials $ 786,600 $ 13,920 F $ 34,200 U
Direct labor $ 318,000 $ 4,200 U $ 15,900 U
Variable manufacturing overhead $ 186,000 $ 4,400 F $ ? U

*Applied to Work in Process during the period.

The company's manufacturing overhead cost is applied to production on the basis of direct labor-hours. All of the materials purchased during the period were used in production. Work in process inventories are insignificant and can be ignored.

Required:

1. How many units were produced last period?

2. How many pounds of direct material were purchased and used in production?

3. What was the actual cost per pound of material? (Round your answer to 2 decimal places.)

4. How many actual direct labor-hours were worked during the period?

5. What was the actual rate paid per direct labor-hour? (Round your answer to 2 decimal places.)

6. How much actual variable manufacturing overhead cost was incurred during the period?
Answer #1

1. No of units produced = Total standard cost/standard cost per unit =786,600/39.33= 20,000 units

2. Pounds of direct material Purchased and used in production:

Material quantity variance = (standard quantity-actual quantity) standard price

                           -34,200 = (20,000*2.30-X) 17.10

                 -34,200 = (46,000-x) 17.10

                          -34,200   = 786,600-17.10X

                           17.10X = 820,800

X (actual quantity purchased and used) = 48,000 pounds

3. Actual cost per pound:

Material price variance = (standard price-actual price) actual quantity

                      13,920 = (17.10-X) 48,000

                      13,920 = 820800-48,000X

                   48,000X     = 806,880

X (actual cost per pound) = 16.81

4.) Calculate actual direct labor hours worked:

Direct labor efficiency variance = (standard hours-actual hours) standard rate

                          -4,200= (20,000-X) 15.90

                             -4,200 = 318,000-15.90X

                                  15.90X     = 322,200

     X (actual direct labor hours)   = 20,264.15 Hours

5. Actual rate per direct labor hour:

Labor rate variance = (standard rate-actual rate) actual hours

                  -48,000     = (15.90-X) 20,264.15

                  -48,000   = 275400-20.26415X

                20.26415X   =323,400

X) (Actual rate per directlabour hour) = 15959.2 per hour

6) Actual variable overhead cost:

Actual variable overhead = standard overhead+Unfavourable variance

                                    = 186000+1943.405

           Unfavorable variance     =20,000- 20,264.15                     

                                                   =-264.15*9.30=-2456.595

                                                   =-2456.595+4400

             Unfavorable variance =1943.405

       Actual variable overhead = 186000+1943.405

Actual variance overhead = 20,543.405